OpenStudy (fuzzyball):

MEDAL AND FAN! Simplify. A. –28xy B. –28 C. 28xy D. 27

OpenStudy (welshfella):

do the multiplying from left to right ( -1/2)*(4x) = -2x now multiply -2x by -2 and so on

OpenStudy (welshfella):

what is -2x * 2 ?

OpenStudy (fuzzyball):

-4

OpenStudy (welshfella):

sorry -2x * - 2 ?

OpenStudy (fuzzyball):

4

OpenStudy (welshfella):

- no youve left out the x its 4x

OpenStudy (welshfella):

now 4x * 7y = ?

OpenStudy (welshfella):

remember x * y = xy

OpenStudy (fuzzyball):

28xy

OpenStudy (welshfella):

right! now multiply by -1 and you have the answer

OpenStudy (fuzzyball):

-28xy

OpenStudy (welshfella):

correct

OpenStudy (fuzzyball):

THANKS SOOO MUCH!!

OpenStudy (welshfella):

yw

OpenStudy (fuzzyball):

can you help me with more?

OpenStudy (welshfella):

no sorry I really have to go right now